please help me in this question....
10^6 -2.5*10^-1*10^-6

Answer by stanbon(75887)

10^6 -2.5*10^-1*10^-6
=10^6 - 2.5*10^-7
=10^6 -(2.5/10^7)
=(10^13 - 2/5)/(10^7)
Cheers,
Stan H.